# Cieran: Designing Sequential Colormaps via In-Situ Active Preference Learning

### 25 Feb 2024 v1

@@ 1. Introduction
Using color in visualization can help intuitively communicate data patterns. We focus on the design of sequential colormaps, which must be orderable, appear smooth and perceptually uniform, and include discriminable colors (Bujack et al., 2018). Beyond these perceptual guidelines, colormaps should also have aesthetic appeal (Moreland, 2009). The aesthetic value of a colormap can depend on color-data semantics (e.g., cool–warm, positive–negative affects) (Bartram et al., 2017), branding, domain conventions, or personal preferences (Palmer et al., 2013). Furthermore, how we see colors applied on a chart will vary with the sizes and shapes of the colored marks (Szafir, 2017), their spatial distribution (Zimnicki et al., 2023), and the chart’s background color (Palmer et al., 2013). These factors and their interactions make designing colormaps challenging for a typical data analyst.
@@
Contributions. Our main technical contributions (§5) include 1. an active learning strategy to induce a personalized model of aesthetic utility for colormaps from few inputs, and 2. a fast algorithm to create new colormaps given the utility model. We validate our technical approaches by conducting a user study (§6) where domain experts ($N=14$) across physical and social sciences took part in an open-ended design task. Our findings show that Cieran can efficiently suggest and create aesthetically pleasing colormaps tailored to individual users in around two minutes of use.
@@ 2.1. Quantifying Colormap Utility
Experts acknowledge that chromatic colormaps should also be aesthetically pleasing (Moreland, 2009). But the aesthetic utility of a colormap will vary with the characteristics of the audience (e.g., personal preference or past experience (Ahmad et al., 2021)), the specific dataset (e.g., semantics (Schloss et al., 2020) and distributions (Zimnicki et al., 2023)), and the problem space (e.g., the target visualization (Stone, 2016) or the data domain (Dasgupta et al., 2018)). For example, how we perceive colors on a chart varies depending on the sizes, the shapes, and the spatial distribution of the colored marks (Szafir, 2017; Zimnicki et al., 2023), as well as the chart’s background color (Palmer et al., 2013).
Consequences of such perceptual dependencies are so prevalent that tools like Tableau [^3] https://www.tableau.com/ automatically modify colormaps for different visualization techniques; for example, by decreasing the color lightness values for area charts and choropleth maps (Brown, 2021). However, such modifications do not take into account dimensions of aesthetics other than the chart technique used, and quantifying the aesthetic value of a colormap with more specificity remains an open problem in visualization (Lau and Moere, 2007; He et al., 2022).
@@ 2.2. Tools for Designing Sequential Colormaps
To help domain experts customize colormaps optimized for their target context, ColorBrewer (Harrower and Brewer, 2003) and matplotlib (Hunter, 2007) and communities like ColourLovers provide expert-designed colormaps to select from. However, sifting through hundreds of recommended colormaps in search of an option with the right look-and-feel can be time-consuming, and default options may not satisfy user needs.
@@
On the other hand, experts have created hundreds of colormaps that tools can leverage as examples. ColorCrafter (Smart et al., 2020) first mines a corpus of 222 colormaps for their structural features and outputs nine model curves that are characteristic of common design practices. Its user can apply a seed color to a model curve and customize it using affine transforms. ColorMoves (Samsel et al., 2018) allows users to mix-and-match segments of example colormaps. While this paradigm may allow for greater expressivity, the tool entrusts users with the combinatorial task of mixing and matching colormap segments.
We aim to reduce the time and expertise needed to both explore a large corpus of colormaps and create new ones when designing a visualization by using active preference learning to model the utility of example colormaps. This model can then be used to rank colormaps and automate the assembly of new colormap structures.
@@ 2.3. Preference-Based Design Optimization
We instead model the aesthetic utility of non-parametric sequential colormaps using a preference-based reward learning (PbRL) algorithm (Cheng et al., 2011). PbRL algorithms are typically used to steer the behavior of virtual and physical robots to complement inverse reinforcement learning (Abbeel and Ng, 2004; Knox and Stone, 2009) or reinforcement learning from human feedback (Akrour et al., 2012; Thomaz and Breazeal, 2008). We hypothesize that they can also guide colormap design as colormaps are continuous and smooth trajectories in a 3D space which a steerable robot might traverse, with the strict constraint that their trajectories monotonically vary in lightness.
Cieran uses this intuition to model the utility of expert-designed colormap trajectories from a small set of pairwise choices to potentially capture a wide range of aesthetic considerations; this utility model can address the needs of visualization designers by rapidly ranking and creating colormaps.
@@ 3. System Objectives
Quantifying the aesthetic value of colormaps remains an open problem in visualization (Lau and Moere, 2007; He et al., 2022). Factors that are difficult to generalize such as personal preferences, data semantics, and data patterns can influence the aesthetic appeal of colormaps. Although galleries like ColorBrewer (Harrower and Brewer, 2003), matplotlib’s colormap library, [^5] https://matplotlib.org/stable/users/explain/colors/colormaps.html and online communities like ColourLovers allow users to explore a range of expert-designed options, manually sifting through colormaps takes significant trial and error. Furthermore, default options limit expressivity. Tools like Photoshop or CCC-Tool can support custom colormap creation, but they require considerable manual tuning, along with the time and expertise to effectively navigate this process. Our goal is to simplify the process of customizing sequential colormaps by reducing the need for user expertise and manual cost.
Our approach builds on the work by Smart et al. (Smart et al., 2020), which mined structural patterns across existing expert-designed colormap curves. By incorporating their colormap corpus within a data analysis software, we aim to enable people to efficiently compare many example colormaps in the context of the data visualization being designed, and in the process learn a model of aesthetic utility for colormaps within the user-specific context. We then aim to use both the learned utility model and the expert examples to create new aesthetically pleasing colormaps to suggest as additional options. Recent enhancements to CCC-Tool (Nardini et al., 2021) include post-processing steps to ensure colormaps meet mathematical formulations of perceptual guidelines outlined by Bujack et al. (Bujack et al., 2018). Their work suggests that tools to create sequential colormaps can and should automatically enforce perceptual guidelines with either constraints or post-processing algorithms.
@@
Figure 3. Cieran is a path planning agent that helps rank and create sequential colormaps with a human in the loop. Cieran first constructs a graphical environment of possible colormap trajectories through a user-specified seed color based on expert demonstrations of colormaps (second subfigure). It then induces a model of aesthetic utility given a small number of pairwise comparison data (the middle figure highlights the user’s selection) collected through a Jupyter Widget, which adaptively and iteratively presents pairs of best candidate colormap examples applied to the user’s target dataset. Using the learned utility model, Cieran will score all expert-designed colormaps (second-to-last subfigure), and create a new quality colormap via path planning (last subfigure).
@@
Figure 3. Cieran is a path planning agent that helps rank and create sequential colormaps with a human in the loop. Cieran first constructs a graphical environment of possible colormap trajectories through a user-specified seed color based on expert demonstrations of colormaps (second subfigure). It then induces a model of aesthetic utility given a small number of pairwise comparison data (the middle figure highlights the user’s selection) collected through a Jupyter Widget, which adaptively and iteratively presents pairs of best candidate colormap examples applied to the user’s target dataset. Using the learned utility model, Cieran will score all expert-designed colormaps (second-to-last subfigure), and create a new quality colormap via path planning (last subfigure).
@@ 5. Cieran: Technical Details
Since colormaps are three-dimensional curves across a colorspace, the problem of finding an aesthetically pleasing colormap is akin to finding a high-utility path within an environment. Cieran executes on this idea by formulating a search space of colormap trajectories, learning a utility model for trajectories from user preferences, and using this model to both rank example colormaps and search for new paths through a colorspace.
@@ 5.1. Technical Overview
Searching for a Novel Colormap. Once the user’s reward model $R(\Upsilon)$ is learned, we could, in principle, use it to score every trajectory in the DAG to find the best novel trajectory (i.e., a combination of piecewise segments from existing colormaps). However, an algorithm that enumerates and scores every trajectory would have exponential time complexity, as it must elicit all possible combinations of $S$ and $A$ (Tarjan, 1978). Given a graph-based environment as well as local action costs, dynamic programming algorithms can achieve path planning to find only the optimal path. However, we define the aesthetic utility of a colormap as a holistic quality (i.e., the harmony of each color to all other colors), not as the sum of utilities across adjacent pairs of colors; such non-additive cost functions preclude the use of combinatorial optimization algorithms. Section 5.4 provides details about an approximate dynamic programming algorithm that can still efficiently search for novel colormaps with high cumulative rewards in our graph-based environment.
@@ 5.2. Colormap Quantization
We first define $(S,A)$, the graph-based environment from which Cieran will sample candidate trajectories through the CIELAB colorspace. This DAG comprises quantized structures from a corpus of 222 expert-designed colormaps from commercial platforms (e.g., Tableau and R) and communities of practice (e.g., ColourLovers) collated in prior work ColorCrafter (Smart et al., 2020). Using this corpus constrains the state and action spaces to those reflecting expert designer practices, and quantizing each colormap allows Cieran to create novel colormaps by joining piecewise segments of existing colormaps.
@@ 5.3. Ranking Expert-Designed Colormaps
For each trajectory $\Upsilon$ through the DAG, we compute a feature vector $\Phi(\Upsilon)$ that encodes its structure according to the parameters in Table 1. Each feature has a corresponding unknown parameter in $\theta$, a weight vector representing the learned preference model, such that:
@@
Algorithm 1 and Figure 1 describe the strategy to actively query pairs of expert-designed colormaps for a user to compare in order to adaptively learn $\theta$.
@@ 5.3.1. Reward Model
Our reward model considers that (1) curves should go through preferred colors (e.g., red) or color families (e.g., warm colors) that a user finds aesthetically pleasing, (2) colormaps should be long enough to be discriminable, but not too long to introduce artifacts like hue banding (Bergman et al., 1995) and (3) that lightness and chroma characteristics are key factors for effective sequential colormaps (Bujack et al., 2018). The reward model has nine path features $\{k_{1},\dots,k_{8}\}$ and $m$ (Table 1) with corresponding unknown weights $\theta_{i}\in[-1,1]$, the preference model learned from user preferences.
@@ 5.5. Implementation
We implement our preference learning algorithm (Algorithm 1) with modifications to the APReL package (Bıyık et al., 2022), which relies on numpy (Harris et al., 2020) and scipy (Virtanen et al., 2020). We implement our adaptive search algorithm (Algorithm 2) with numpy and networkx (Hagberg et al., 2008). The processing pipeline to interpolate a B-spline curve from each trajectory and sample $\Delta E_{2000}$-equidistant colors is implemented using coloraide (Muse, 2023). Each colormap is an instance of matplotlib’s ListedColormap class (Hunter, 2007). The interface is implemented using ipywidgets (Jupyter Widgets Team, 2023).
@@ 6.1.4. Quantitative data collection:
To reduce potential confounds from color pickers, participants specified a seed color from the Tableau10 color palette (Stone, 2016), an expert-designed palette of ten discrete colors designed for categorical visualization. The first phase used the color in Tableau10 that was closest to the participant’s favorite color, but participants were free to choose any color, including a previously used color, in other phases. For each pairwise comparison, the participant could indicate their preference for left (1), right (2), or indifferent (0) input using a textbox. They were instructed to choose 1 or 2 if there was even the slightest preference for one of the colormaps or if they simply disliked one more than the other. These comparisons served as inputs to model each participant’s preferences using the algorithm described in Section 1.
@@ 6.2.1. Quantitative Results
We report inferential statistics, means, and standard errors (means $\pm$ standard errors) for the differences in ranked-choice votes (or, worth [^6] https://cran.rstudio.com/web/packages/PlackettLuce/vignettes/Overview.html) across Cieran’s new and example colormaps.
@@ 8. Future Work
Our system is limited by the number of expert-designed colormaps in the corpus. Furthermore, displays have a limited color gamut. When colormaps are rotated and/or translated to fit a seed color, Cieran discards some colormap trajectories because they run outside of the gamut and produce colors that many monitors cannot display. This gamut constraint motivated us to use the Tableau10 color palette when limiting the selection of seed colors in the user study since colormaps aligned to these colors were well within the gamut boundaries. Although Tableau10 was designed by a visualization expert, we plan to extend the colormap corpus to accommodate as many seed colors as possible. Analysts’ learned preference models can be re-used to sort a growing corpus with more options. We also plan to compile a recommended list of ‘good’ seed colors.
@@
Lastly, while Cieran can rank and create variants of both single-hue and multi-hue colormaps effectively, prior studies have examined performance trade-offs across these colormap categories for specific data domain tasks, such as spatial data analysis (Dasgupta et al., 2018) and model checking (Reda and Szafir, 2021). The results illustrate that each colormap type may have its advantages depending on the domain task. Like many visualization construction tools, Cieran assumes that its users have an understanding of their target context, and its algorithms are not guardrails to prevent the construction of potentially inadequate visualizations: it simply acts on user inputs to help users quickly create visualizations that meet universal perceptual guidelines. Future iterations of Cieran could provide additional scenario-driven support, such as automatically adjusting hue variance based on a set of target tasks or privileging hues that align with domain conventions in addition to seed colors.

### 29 Feb 2024 v2

@@ 1. Introduction
When visualizing data, analysts frequently need to decide what colors to use in a chart. We focus on the design of sequential colormaps, which must be orderable, appear smooth, include discriminable colors, and adhere to perceptual uniformity (Bujack et al., 2018). Beyond these perceptual guidelines, colormaps should also have aesthetic appeal (Moreland, 2009), which can depend on color-data semantics (e.g., cool–warm, positive–negative affects) (Bartram et al., 2017), branding, domain conventions, or personal preferences (Palmer et al., 2013). Furthermore, how we see colors applied on a chart will vary with the sizes and shapes of the colored marks (Szafir, 2017), their spatial distribution (Zimnicki et al., 2023), and the chart’s background color (Palmer et al., 2013). These factors and their interactions make colormap design challenging for a typical data analyst.
@@
Contributions. Our main technical contributions (§5) include 1. an active learning strategy to induce a personalized model of aesthetic utility for colormaps from few inputs, and 2. a fast algorithm to create new colormaps given the utility model. We validate our technical approaches by conducting a user study (§6) where domain experts ($N=12$) across physical and social sciences took part in an open-ended design task. Our findings show that Cieran can efficiently suggest and create aesthetically pleasing colormaps tailored to individual users in around two minutes of use.
@@ 2.1. Quantifying Colormap Utility
Experts acknowledge that chromatic colormaps should also be aesthetically pleasing (Moreland, 2009). But the aesthetic utility of a colormap will vary with the characteristics of the audience (e.g., personal preference or past experience (Ahmad et al., 2021)), the specific dataset (e.g., semantics (Schloss et al., 2020) and distributions (Zimnicki et al., 2023)), and the problem space (e.g., the target visualization (Szafir, 2017) or the data domain (Dasgupta et al., 2018)). For example, how we see colors on a chart varies depending on the sizes, the shapes, and the spatial distribution of the colored marks (Szafir, 2017; Zimnicki et al., 2023).
Consequences of these factors and their interactions are so prevalent that tools like Tableau [^3] https://www.tableau.com/ automatically modify colormaps for different visualization techniques; for example, by decreasing the color lightness values for area charts and choropleth maps (Brown, 2021). However, such modifications do not take into account dimensions of aesthetics other than the chart technique used, and quantifying the aesthetic value of a colormap with more specificity remains an open problem in visualization (Lau and Moere, 2007; He et al., 2022).
@@ 2.2. Tools for Designing Sequential Colormaps
To help domain experts customize colormaps for their target context, ColorBrewer (Harrower and Brewer, 2003), matplotlib (Hunter, 2007) and online communities provide galleries of expert-designed colormaps. However, searching across hundreds of colormaps for the right look-and-feel can be time-consuming, and default options may not satisfy user needs.
@@
Instead, systems can directly leverage the hundreds of expert-designed colormaps as examples. ColorCrafter (Smart et al., 2020) first mines a corpus of 222 colormaps for their structural features and outputs nine model curves that are characteristic of common design practices. People can apply a seed color to one of the model curves and customize it using affine transforms. ColorMoves (Samsel et al., 2018) allows users to mix-and-match segments of example colormaps. While such a paradigm may be expressive, the tool entrusts users with the combinatorial task of mixing and matching colormap segments.
In our work, we aim to build a tool that automatically suggests and creates colormaps from examples, in a manner that is driven by users’ aesthetic preferences. We hypothesize that active preference learning can be used to rank colormaps according to aesthetic utility and automate the assembly of new colormap structures.
@@ 2.3. Preference-Based Design Optimization
We instead model the aesthetic utility of non-parametric sequential colormaps using a preference-based reward learning (PbRL) algorithm (Cheng et al., 2011). PbRL algorithms are typically used to steer the behavior of virtual and physical robots to complement inverse reinforcement learning (Abbeel and Ng, 2004; Knox and Stone, 2009) or reinforcement learning from human feedback (Akrour et al., 2012; Thomaz and Breazeal, 2008). We hypothesize they can also guide the design of colormaps, which are continuous and smooth trajectories in a 3D space that a steerable robot might traverse.
Cieran uses this intuition to model the aesthetic utility of colormap curves from a few pairwise comparisons to potentially capture a wide range of aesthetic considerations, and in turn, optimize the design of sequential colormaps.
@@ 3. System Objectives
Quantifying the aesthetic value of colormaps remains an open problem in visualization (Lau and Moere, 2007; He et al., 2022). Factors that are difficult to generalize such as personal preferences, data semantics, and data patterns can influence the aesthetic appeal of colormaps. Although galleries like ColorBrewer, matplotlib’s colormap library, and online communities like ColourLovers allow users to explore a range of expert-designed options, manually sifting through colormaps takes significant trial and error. Furthermore, default options limit expressivity. Tools like Photoshop or CCC-Tool can support custom colormap creation, but they require considerable manual tuning, along with the time and expertise to effectively navigate this process. Our goal is to simplify the process of customizing sequential colormaps by reducing the need for user expertise and manual cost.
Our approach builds on the work by Smart et al. (Smart et al., 2020), which mined structural patterns across existing expert-designed colormap curves. By incorporating their colormap corpus within a data analysis software, we aim to enable people to efficiently rank many example colormaps in the context of the data visualization being designed, and in the process learn a model of aesthetic utility for colormaps within the user-specific context. We then aim to use both the learned utility model and the expert examples to create new aesthetically pleasing colormaps to suggest as additional options. Recent enhancements to CCC-Tool (Nardini et al., 2021) include post-processing steps to ensure colormaps meet mathematical formulations of perceptual guidelines outlined by Bujack et al. (Bujack et al., 2018). Their work suggests that tools to create sequential colormaps can and should automatically enforce perceptual guidelines with either constraints or post-processing algorithms.
@@
Figure 3. Cieran is a path planning agent that helps rank and create sequential colormaps with a human in the loop. Cieran first constructs a graphical environment of possible colormap trajectories through a user-specified seed color based on expert demonstrations of colormaps (second subfigure). It then induces a model of aesthetic utility given a small number of pairwise comparison data (the third subfigure highlights the user’s selection) collected through a Jupyter Widget, which adaptively and iteratively presents pairs of best candidate colormap examples applied to the user’s target dataset. Using the learned utility model, Cieran will score all expert-designed colormaps (fourth subfigure), and create a new quality colormap via path planning (fifth subfigure).
@@
Figure 3. Cieran is a path planning agent that helps rank and create sequential colormaps with a human in the loop. Cieran first constructs a graphical environment of possible colormap trajectories through a user-specified seed color based on expert demonstrations of colormaps (second subfigure). It then induces a model of aesthetic utility given a small number of pairwise comparison data (the third subfigure highlights the user’s selection) collected through a Jupyter Widget, which adaptively and iteratively presents pairs of best candidate colormap examples applied to the user’s target dataset. Using the learned utility model, Cieran will score all expert-designed colormaps (fourth subfigure), and create a new quality colormap via path planning (fifth subfigure).
@@ 5. Cieran: Technical Details
Since colormaps are three-dimensional curves across a colorspace, the problem of finding an aesthetically pleasing colormap is akin to finding a high-utility path within an environment. Cieran executes on this idea by formulating a search space of colormap trajectories, learning a utility model for trajectories from user preferences, and using this model to both rank example colormaps and search for new paths through a colorspace (Figure 3).
@@ 5.1. Technical Overview
Searching for a Novel Colormap. Once $R(\Upsilon)$ is learned, we could, in principle, use it to score every trajectory in the DAG to find the best novel trajectory (i.e., a combination of piecewise segments from existing colormaps). However, an algorithm that enumerates and scores all possible combinations of $S$ and $A$ would have exponential time complexity (Tarjan, 1978). Given a graph, dynamic programming algorithms can achieve path planning to find only the optimal path. However, we define the aesthetic utility of a colormap as a holistic quality (i.e., the harmony of each color to all other colors), not as the sum of utilities across locally adjacent color pairs; such non-additive cost functions preclude the use of combinatorial optimization. Section 5.4 details an approximate dynamic programming algorithm that can still efficiently search for novel colormaps with high cumulative rewards in our graph-based environment.
@@ 5.2. Colormap Quantization
We first define $(S,A)$, the graph-based environment from which Cieran will sample candidate trajectories through the CIELAB colorspace (Figure 3, first subfigure). This DAG comprises quantized structures from a corpus of 222 expert-designed colormaps from commercial platforms (e.g., Tableau and R) and communities of practice (e.g., ColourLovers) collated in prior work ColorCrafter (Smart et al., 2020). Using this corpus constrains the state and action spaces to those reflecting expert designer practices, and quantizing each colormap allows Cieran to create novel colormaps by joining piecewise segments of existing colormaps.
@@ 5.3. Ranking Expert-Designed Colormaps
For each trajectory $\Upsilon$ through the DAG, we can define a feature vector $\Phi$ that describes its characteristics. Then $R$, the utility of a trajectory, can be defined as a linear combination these features, such that:
@@
Algorithm 1 and Figure 5 describe the strategy to actively learn $\theta$ from pairwise colormap preferences, which requires the specification of reward features, a belief model, and an acquisition model.
@@ 5.3.1. Reward Model
Our reward model considers that (1) curves should go through preferred colors (e.g., red) or color families (e.g., warm colors) that a user finds aesthetically pleasing, (2) colormaps should be long enough to be discriminable, but not too long to introduce artifacts like hue banding (Bergman et al., 1995) and (3) that lightness and chroma characteristics are key factors for effective sequential colormaps (Bujack et al., 2018). The reward model has nine path features $\{k_{1},\dots,k_{8}\}$ and $m$ (Table 1) with corresponding unknown weights $\theta_{i}\in[-1,1]$, the preference model learned from user choices.
@@ 5.5. Implementation
We implement Algorithm 1 with modifications to the APReL package (Bıyık et al., 2022), which uses numpy (Harris et al., 2020) and scipy (Virtanen et al., 2020). We implement Algorithm 2 with numpy and networkx (Hagberg et al., 2008). B-spline interpolation of curves and sampling $\Delta E_{2000}$-equidistant colors utilize coloraide (Muse, 2023). Each colormap is a matplotlib.ListedColormap (Hunter, 2007) class instance. We implement the interface using ipywidgets (Jupyter Widgets Team, 2023).
@@ 6.1.4. Quantitative data collection:
To reduce potential confounds from color pickers, participants specified a seed color from the Tableau10 color palette (Stone, 2016), an expert-designed palette of ten discrete colors designed for categorical visualization. The first phase used the color in Tableau10 that was closest to the participant’s favorite color, but participants were free to choose any color, including a previously used color, in other phases. For each pairwise comparison, the participant could indicate their preference for left (1), right (2), or indifferent (0) input using a textbox. They were instructed to choose 1 or 2 if there was even the slightest preference for one of the colormaps or if they simply disliked one more than the other.
@@ 6.2.1. Quantitative Results
We report inferential statistics, means, and standard errors (means $\pm$ standard errors) for the differences in ranked-choice votes (or, worth [^5] https://cran.rstudio.com/web/packages/PlackettLuce/vignettes/Overview.html) across Cieran’s new and example colormaps (Figure 7).
@@ 8. Future Work
Our system is limited by the number of expert-designed colormaps in the corpus. Furthermore, displays have a limited color gamut. When colormaps are rotated and/or translated to fit a seed color, Cieran discards some colormap trajectories because they run outside of the gamut and produce colors that many monitors cannot display. This gamut constraint motivated us to use the Tableau10 color palette when limiting the selection of seed colors in the user study since colormaps aligned to these colors were well within the gamut boundaries. Although Tableau10 was designed by a visualization expert, we plan to extend the colormap corpus to accommodate as many seed colors as possible. We also plan to compile a recommended list of ‘good’ seed colors.
@@
Lastly, while Cieran can rank and create variants of both single-hue and multi-hue colormaps effectively, users should be aware of performance trade-offs between these colormap categories for specific data domain tasks, such as spatial data analysis (Dasgupta et al., 2018) and model checking (Reda and Szafir, 2021). Like many visualization construction tools, Cieran assumes that its users have an understanding of their target context, and its algorithms are not guardrails to prevent the construction of potentially inadequate visualizations: it simply acts on user inputs to help users quickly create visualizations that meet universal perceptual guidelines. Future iterations of Cieran could provide additional scenario-driven support, such as automatically adjusting hue variance based on a set of target tasks or privileging hues that align with domain conventions in addition to seed colors.